What is the volume of a tantalum Atom in microliters
1
Expert's answer
2020-08-27T05:48:11-0400

Volume of tantalum atom is "1.2\times 10^{-23} \ cm^3=1.2\times 10^{-23} \ ml"

And "1" milliliter = "10^{3}\" microliter "(\mu L)"

So, Volume of tantalum atom is "1.2\times 10^{-23} \ \times 10^3=1.2\times 10^{-20}\ \mu L"